# Main-belt and Trojan Asteroid Phase Curves from the ATLAS Survey

### v1

@@ 3.2 Comparison with previous studies
Figure 10: Histograms showing the difference in $H$ (upper panel) and $G_{12}$ (lower panel) of our phase curve fits in the P16 system, cross-matched with the results of Mahlke et al. (2021) who also used ATLAS data. The median of each distribution is indicated by the red vertical line with values of -0.01 and -0.06 magnitudes for the upper and lower panels respectively. The vertical black lines mark the zero value of the $x$-axis.
@@
Figure 12: Histogram comparing the ATLAS $H_{c}-H_{o}$ colour distribution of objects in this work cross-matched with the previous colours ($c-o$) published by McNeill et al. (2021), who also investigated the properties of the Jupiter Trojans using ATLAS data. The median value of the distribution is -0.008 magnitudes, the vertical black line indicates the zero value of the $x$-axis.
@@ 4.2 Asteroid families in the SAP database
Figure 13: Proper orbital element distributions of MBAs in the SAP database (elements from astdys). The proper semimajor axis is plotted against the sine of proper inclination (upper panel) and proper eccentricity (lower panel). The 63,197 non-family members in the dataset are displayed as a two dimensional histogram in greyscale. The 25,479 asteroids that are a member of a family (as classified by astdys) are colour-coded by $G$ with a diverging colour map, where the colour transition is centred on the median value of $G$. The same colour scheme is used for both plots and objects with phase parameters outside the $1-99\%$ quantile of the family member $G$ distribution were excluded from the plot in order to avoid stretching of the colour map by the small number of more extreme values. As such we only display asteroid family members with $0.600.03<G<$ 0.60 (where the median $G$ of this sample is 0.26). Some of the larger asteroid families are labelled by black crosses at their mean proper elements (Milani et al., 2014).
Figure 14: Here we show the astdys proper orbital element distribution of Nysa-Polana (N-P) asteroid family members in the SAP dataset, in proper eccentricity vs. proper semimajor axis space (left panel). We indicate the proper elements of the main asteroids of the complex (Nysa, Polana and Hertha) with shaped markers. Every other family member is colour coded according to a two dimensional colour map, $H_{c}-H_{o}$ colour against phase parameter $G$. This colour map is displayed in the right-hand panel, where we have also marked the locations of the SAP Nysa-Polana family members in this parameter space (black circles). The black contour indicates the 10% level relative to the maximum count of the underlying two dimensional histogram of the distribution. We have also indicated the median $H_{c}-H_{o}$ and $G$ of the S and C-type asteroids in the full SAP database (red cross and plus respectively, values from figure 4).
@@ 4.3 Jupiter Trojan asteroids in the SAP database
Here we consider the phase curve properties of the Jupiter Trojans present in our SAP dataset. This contains 740 Jupiter Trojans with high quality $HG$ fits in the $o$-filter, of which 570 also have good fits in the $c$-filter suitable for calculating a $H_{c}-H_{o}$ colour (section 2.6). 103/740 of these Jupiter Trojans display strong apparition effects and our sample contains 1.98 more Jupiter Trojans at L4 than L5. This is similar to the L4/L5 number ratio for observed Jupiter Trojans (for counts of Jupiter Trojans recorded by the MPC, Li et al., 2022) but is larger than the number ratio of $\sim 1.4$ when the underlying size-frequency distributions are accounted for (Grav et al., 2011; Uehata et al., 2022). We compare our distributions in absolute magnitude, $H$, of these two groups (figure 15, left panels) using a two sample Kolmogorov-Smirnov (KS) test. In a two sample KS test the parameter $D$ describes the maximum distance (at a given value) between the cumulative distributions of the two datasets being tested. The KS statistic is calculated with a $p$-value which indicates the probability that the null hypothesis is true, i.e. that both datasets were drawn from the same underlying distribution. Conventionally a result is statistically significant when $p<0.05$ (see also equation 10). For the L4/L5 $H$ distributions this test results in a KS statistic of $D=0.2$ (with a $p$-value of $3\times 10^{-6}$) from which we reject the null hypothesis that the L4 and L5 absolute magnitudes are drawn from the same distribution. There is no strong evidence in the literature that the L4, L5 groups should have different size distributions (e.g. Uehata et al., 2022). In this case we determined that the differences in $H$ were introduced into our study through the initial watchlist of asteroids monitored by rockAtlas. There is no significant difference in the absolute magnitude distribution for the L4/L5 Trojans when using $H_{V}$ from the full astorb dataset. However when we compare the astorb $H_{V}$ of only the L4/L5 Trojans which were selected for the rockAtlas watchlist, a KS test displays statistically significant differences similar to those seen in figure 15 (left). Therefore we can rule out observational bias between the L4 and L5 groups in the ATLAS survey as the main cause of this effect, and attribute it instead to selection bias.
@@
Figure 15: Histogram distributions of $H$, $G$ and $H_{c}-H_{o}$ colour, in the B89 system, for the L4 and L5 Jupiter Trojans in the SAP dataset (top row, left, middle and right panels respectively). The distributions for the L4 and L5 groups are shown in blue and orange respectively; the median of each distribution is indicated by the vertical coloured line. The cumulative distribution functions (CDFs) of each parameter are shown in the bottom row, allowing us to compare the distributions via the two-sample Kolmogorov-Smirnov test. The KS statistic $D$ is determined from the maximum difference between the CDFs at a given value and when this is sufficiently large (at a level determined by the $p$-value) we can reject the null hypothesis that the two datasets are drawn from the same underlying distribution.
@@
Figure 16: Scatter plot of phase parameter $G$ from the SAP dataset against the ZTF $g-r$ colours from Schemel & Brown (2021), for objects matched between the two datasets (upper panel). Similar to their analysis we only include objects from both datasets with lower than median uncertainties (compare to their figure 6). The red vertical line corresponds to the colour boundary at $g-r=0.565$ which divides the “red” and “less-red” Jupiter Trojan compositional classes The lower panel shows the histogram distributions of ATLAS $G$ for the red and less-red Jupiter Trojans from the panel above.
@@ 4.4.2 Catalogue contamination by poor quality fits
One of the complex aspects of this work was the production of a final high quality phase curve dataset. We fitted all available apparitions and then ideally one should be able to use the uncertainties of the fitted phase curve parameters to assess the quality of the dataset. However as discussed previously, we do not account for rotational brightness variation and therefore the formal uncertainties produced by the Levenberg-Marquadt algorithm do not reflect the true range of possible values for the phase curve parameters. We therefore had to develop a series of cuts based on a variety of metrics which we tried to choose in a relatively agnostic way so as to not overly influence the final results (section 2.5 and appendix A). We did this by considering especially the phase curves of objects with unusually high values of $G$ and $\sigma_{G}$ and adjusting our metrics such that these truly unphysical fits were rejected. We acknowledge that this approach relies on somewhat subjective decisions. A more realistic estimate of the phase parameter uncertainties could have been obtained by methods in which rotational variation is accounted for (e.g. SB21, ). However, this would add further computational cost to calculating phase parameters for each apparition of a large number of objects. We did consider “quick” solutions to deal with rotational effects, e.g. binning the observed magnitude over some timescale or by fitting random samples of the observations, however we found that these methods did not significantly improve the quality of the final dataset.
@@ 4.4.3 Phase curve wavelength dependence
We determined the phase parameter ($G$, $G_{12}$) for each object from its primary apparition in the $o$-filter, and kept this phase parameter fixed when fitting all other apparitions. There is evidence that phase curve properties vary with wavelength in a “phase-coloring” effect (Alvarez-Candal et al., 2022a; Ďurech et al., 2020). However in the ATLAS dataset the coverage in the $c$-filter is much sparser compared to the $o$-filter. In most cases we lacked sufficient data to obtain an independent phase curve for the $c$-filter data. In order to obtain usable colour information we were forced to assume the slope parameters to be wavelength independent in order to find $H_{o}$ and $H_{c}$. This is a reasonable assumption given the accuracy of ATLAS photometry and not accounting for rotational variation. These effects probably dominate over any changes in phase parameter due to the wavelength of observation.
@@ 4.4.4 Rotational lightcurve effects
While developing our methodology we attempted several simple methods to incorporate rotational brightness variations into our phase curve fits. For example we considered a Monte Carlo approach whereby each observation was re-sampled based on its measurement uncertainty and phase curves were repeatedly fit (similar to McN21, ). We also attempted to average out rotational variation by binning reduced magnitudes nightly when fitting the phase curves. However, we found that these methods did not adequately account for rotational variations in the lightcurves. When resampling observations for objects displaying clear lightcurve variations the photometric uncertainty is generally small compared to the amplitude and so the fitted phase curve parameters and their uncertainties did not differ significantly from our original method. Similarly, binning the data did not have much effect on the final parameters. Furthermore, binning observations of objects with a wide possible range of rotational periods (minutes to days) compared to the ATLAS observational cadence would likely introduce a number of additional biases to our analysis. A more robust treatment would account for the unknown rotational phase and amplitude of each observation and so the true uncertainties of the phase curve parameters would be much wider but reflect better the true distribution of possible parameters (e.g. Waszczak et al. (2015); Alvarez-Candal et al. (2022b), SB21). However such methods are generally more computationally intensive to implement, therefore in this work we stuck with a simple phase curve fitting method, neglecting rotational effects as described in section 2.4. This allowed us to consider a sizeable sample of asteroids, and their individual apparitions, in a timely manner. As such, the formal fit uncertainties produced in this work should be used to compare relative goodness of fit within the dataset; they do not necessarily reflect the true uncertainty in the fitted phase curve parameters if rotational effects are strong.
@@ 4.5 Applicability to future surveys
In this work we have demonstrated that sparse wide field survey data can be exploited to investigate and broadly compare the properties of different asteroid populations. In particular, by considering the phase parameters and surface colours derived in this work one may compare and contrast different dynamical groups such as the asteroid families and Jupiter Trojans. However as discussed above, in order to make more precise assessments one must take into account the rotational variation introduced by the asteroid lightcurve, either in a statistical sense (e.g. Waszczak et al. (2015); Alvarez-Candal et al. (2022b), SB21) or with more detailed lightcurve shape modelling (e.g. Ďurech et al., 2020). Despite our simple phase curve model, we have built upon previous phase curve studies by considering the phase curve properties of individual apparitions for a large sample of over $100,000$ asteroids in the Main Asteroid Belt to Jupiter Trojan region, with data spanning a significant temporal baseline since 2017. By doing so we have been able to identify objects that are undergoing changes in observed absolute magnitude between apparitions. By excluding these objects from our science analysis of absolute magnitudes and colours we have improved the purity of our sample.
@@
The presence of the aforementioned apparition effects must be accounted for when analysing the results of wide field surveys expected to observe large numbers of objects over a long baseline. In particular, LSST will discover approximately $6\times 10^{6}$ MBAs over 10 years, with 100s of observations for most objects. In addition to moving object discovery, LSST will provide automatically derived properties of observed asteroids such as lightcurves, phase curves and colours (LSST Science Collaboration et al., 2009). Given the volume and depth of the survey it will be difficult to followup the expected large numbers of faint asteroids, therefore LSST will likely be the only source of data for many of these objects. Considering the apparition effect is therefore essential in order to determine accurate absolute magnitudes, which directly effects studies regarding the size distribution of small Solar System bodies, such as the formation and evolution of planetesimals, properties of asteroid families, planetary defence, etc.
Furthermore, our approach which has involved detailed analysis of phase curve residuals offers a framework to monitor for unusual observations that might indicate an object of interest. We considered a range of statistics for the phase curve residuals, and in the vast majority of cases any rejected apparitions were due to poor observational coverage or low quality photometry. However, certain physical processes may also lead to deviations from the expected phase curve behaviour, e.g. rapid changes in viewing geometry (Jackson et al., 2022) or sudden increases in brightness due to the onset of activity (Jewitt & Hsieh, 2022). For example, asteroid 2008 GO98 was found to be active in July 2017 (Kokhirova et al., 2021). Our methodology produced a most unusual phase curve for this object, with an extreme value of $G=1675$ which was determined from an apparition spanning from 8th June until 9th December 2017. Future work in this area could involve identifying these outliers in phase curve parameter space and using the ATLAS forced photometry server to get cutouts to search for activity. Continuous updates of derived physical properties (such as absolute magnitude, phase curve and colours) as new observations are made will be an essential part of large surveys such as LSST. Therefore learning how to identify truly unusual objects in this process will greatly advance the study of relatively rare objects such as the active asteroids.
@@ 5 Conclusions
We have compiled a database of ATLAS asteroid photometry by matching the ephemerides of 4.5×$\times 10^{5}$ asteroids with accurate orbital elements to the detections produced by the ATLAS dophot pipeline, using rockAtlas (Young, 2023). We fitted phase curves in the $HG$, $H_{12}G_{12}$ models (Bowell et al., 1989; Penttilä et al., 2016) to the apparition in the $o$-filter with the best combination of number of observations and phase angle coverage. We then fitted all other apparitions for $H$ with these derived phase parameters fixed. This allowed us to assess any changes in $H$ during the survey and also to determine the $H_{c}-H_{o}$ surface colour of many asteroids. We used various metrics to select 106,055 asteroids with well defined phase curve fits for the final Selected ATLAS Phase-curves (SAP) dataset, which is available online.
@@
In this work we chose a phase curve model without rotational lightcurve variations to better investigate individual apparitions of a large number of asteroids. However, we recommend that future studies account for rotational effects, e.g. via likelihood models including rotational variation (for sparse data e.g. Schemel & Brown, 2021) and/or full CLI techniques (for objects with additional lightcurve rotation periods e.g. Ďurech et al., 2020). Accurate phase curve modelling with realistic uncertainties on absolute magnitude and phase curve parameters are essential when comparing the optical properties of different small body populations. This will be particularly relevant for large, sparse surveys such as LSST which will discover and characterise approximately $\times 10^{6}$ asteroids over its 10 year baseline (LSST Science Collaboration et al., 2009). Many LSST objects will be too faint and/or numerous for dedicated followup therefore extracting accurate parameters from sparse data alone, and quickly highlighting interesting outliers, is a pressing issue.
[…]

### v2

@@ 3.2 Comparison with previous studies
Figure 10: Histograms showing the difference in $H$ (upper panel) and $G_{12}$ (lower panel) of our phase curve fits in the P16 system, cross-matched with the results of Mahlke et al. (2021) who also used ATLAS data. The median of each distribution is indicated by the vertical dashed line with values of -0.01 and -0.06 magnitudes for the upper and lower panels respectively. The vertical solid lines mark the zero value of the $x$-axis.
@@
Figure 12: Histogram comparing the ATLAS $H_{c}-H_{o}$ colour distribution of objects in this work cross-matched with the previous colours ($c-o$) published by McNeill et al. (2021), who also investigated the properties of the Jupiter Trojans using ATLAS data. The median value of the distribution is -0.008 magnitudes (vertical dashed line), the vertical solid line indicates the zero value of the $x$-axis.
@@ 4.2 Asteroid families in the SAP database
Figure 13: Proper orbital element distributions of MBAs in the SAP database (elements from astdys). The proper semimajor axis is plotted against the sine of proper inclination (upper panel) and proper eccentricity (lower panel). The 63,197 non-family members in the dataset are displayed as a two dimensional histogram in greyscale. The 25,479 asteroids that are a member of a family (as classified by astdys) are colour-coded by $G$ with a diverging colour map, where the colour transition is centred on the median value of $G$. The same colour scheme is used for both plots and objects with phase parameters outside the $1-99\%$ quantile of the family member $G$ distribution were excluded from the plot in order to avoid stretching of the colour map by the small number of more extreme values. As such we only display asteroid family members with $0.600.03<G<$ 0.60 (where the median $G$ of this sample is 0.26). Some of the larger asteroid families are labelled by crosses at their mean proper elements (Milani et al., 2014).
Figure 14: Here we show the astdys proper orbital element distribution of Nysa-Polana (N-P) asteroid family members in the SAP dataset, in proper eccentricity vs. proper semimajor axis space (left panel). We indicate the proper elements of the main asteroids of the complex (Nysa, Polana and Hertha) with shaped markers. Every other family member is colour coded according to a two dimensional colour map, $H_{c}-H_{o}$ colour against phase parameter $G$. This colour map is displayed in the right-hand panel, where we have also marked the locations of the SAP Nysa-Polana family members in this parameter space (circular markers). The contour indicates the 10% level relative to the maximum count of the underlying two dimensional histogram of the distribution. We have also indicated the median $H_{c}-H_{o}$ and $G$ of the S and C-type asteroids in the full SAP database (cross and plus markers respectively, values from figure 4).
@@ 4.3 Jupiter Trojan asteroids in the SAP database
Here we consider the phase curve properties of the Jupiter Trojans present in our SAP dataset. This contains 740 Jupiter Trojans with high quality $HG$ fits in the $o$-filter, of which 570 also have good fits in the $c$-filter suitable for calculating a $H_{c}-H_{o}$ colour (section 2.6). 103/740 of these Jupiter Trojans display strong apparition effects and our sample contains 1.98 more Jupiter Trojans at L4 than L5. This is similar to the L4/L5 number ratio for observed Jupiter Trojans (for counts of Jupiter Trojans recorded by the MPC, Li et al., 2023) but is larger than the number ratio of $\sim 1.4$ when the underlying size-frequency distributions are accounted for (Grav et al., 2011; Uehata et al., 2022). We compare our distributions in absolute magnitude, $H$, of these two groups (figure 15, left panels) using a two sample Kolmogorov-Smirnov (KS) test. In a two sample KS test the parameter $D$ describes the maximum distance (at a given value) between the cumulative distributions of the two datasets being tested. The KS statistic is calculated with a $p$-value which indicates the probability that the null hypothesis is true, i.e. that both datasets were drawn from the same underlying distribution. Conventionally a result is statistically significant when $p<0.05$ (see also equation 10). For the L4/L5 $H$ distributions this test results in a KS statistic of $D=0.2$ (with a $p$-value of $3\times 10^{-6}$) from which we reject the null hypothesis that the L4 and L5 absolute magnitudes are drawn from the same distribution. There is no strong evidence in the literature that the L4, L5 groups should have different size distributions (e.g. Uehata et al., 2022). In this case we determined that the differences in $H$ were introduced into our study through the initial watchlist of asteroids monitored by rockAtlas. There is no significant difference in the absolute magnitude distribution for the L4/L5 Trojans when using $H_{V}$ from the full astorb dataset. However when we compare the astorb $H_{V}$ of only the L4/L5 Trojans which were selected for the rockAtlas watchlist, a KS test displays statistically significant differences similar to those seen in figure 15 (left). Therefore we can rule out observational bias between the L4 and L5 groups in the ATLAS survey as the main cause of this effect, and attribute it instead to selection bias.
@@
Figure 15: Histogram distributions of $H$, $G$ and $H_{c}-H_{o}$ colour, in the B89 system, for the L4 and L5 Jupiter Trojans in the SAP dataset (top row, left, middle and right panels respectively). The distributions for the L4 and L5 groups are shown as coloured histograms; the median of each distribution is indicated by the matching colour vertical line. The cumulative distribution functions (CDFs) of each parameter are shown in the bottom row, allowing us to compare the distributions via the two-sample Kolmogorov-Smirnov test. The KS statistic $D$ is determined from the maximum difference between the CDFs at a given value and when this is sufficiently large (at a level determined by the $p$-value) we can reject the null hypothesis that the two datasets are drawn from the same underlying distribution.
@@
Figure 16: Scatter plot of phase parameter $G$ from the SAP dataset against the ZTF $g-r$ colours from Schemel & Brown (2021), for objects matched between the two datasets (upper panel). Similar to their analysis we only include objects from both datasets with lower than median uncertainties (compare to their figure 6). The vertical line corresponds to the colour boundary at $g-r=0.565$ which divides the “red” and “less-red” Jupiter Trojan compositional classes The lower panel shows the histogram distributions of ATLAS $G$ for the red and less-red Jupiter Trojans from the panel above.
@@ 4.4.2 Catalogue contamination by poor quality fits
One of the complex aspects of this work was the production of a final high quality phase curve dataset. We fitted all available apparitions and then ideally one should be able to use the uncertainties of the fitted phase curve parameters to assess the quality of the dataset. However as discussed previously, we do not account for rotational brightness variation and therefore the formal uncertainties produced by the Levenberg-Marquardt algorithm do not reflect the true range of possible values for the phase curve parameters. We therefore had to develop a series of cuts based on a variety of metrics which we tried to choose in a relatively agnostic way so as to not overly influence the final results (section 2.5 and appendix A). We did this by considering especially the phase curves of objects with unusually high values of $G$ and $\sigma_{G}$ and adjusting our metrics such that these truly unphysical fits were rejected. We acknowledge that this approach relies on somewhat subjective decisions. A more realistic estimate of the phase parameter uncertainties could have been obtained by methods in which rotational variation is accounted for (e.g. SB21, ). However, this would add further computational cost to calculating phase parameters for each apparition of a large number of objects. We did consider “quick” solutions to deal with rotational effects, e.g. binning the observed magnitude over some timescale or by fitting random samples of the observations, however we found that these methods did not significantly improve the quality of the final dataset.
@@ 4.4.3 Phase curve wavelength dependence
We determined the phase parameter ($G$, $G_{12}$) for each object from its primary apparition in the $o$-filter, and kept this phase parameter fixed when fitting all other apparitions. There is evidence that phase curve properties vary with wavelength in a “phase-coloring” effect (Alvarez-Candal et al., 2022b; Ďurech et al., 2020). However in the ATLAS dataset the coverage in the $c$-filter is much sparser compared to the $o$-filter. In most cases we lacked sufficient data to obtain an independent phase curve for the $c$-filter data. In order to obtain usable colour information we were forced to assume the slope parameters to be wavelength independent in order to find $H_{o}$ and $H_{c}$. This is a reasonable assumption given the accuracy of ATLAS photometry and not accounting for rotational variation. These effects probably dominate over any changes in phase parameter due to the wavelength of observation.
@@ 4.4.4 Rotational lightcurve effects
While developing our methodology we attempted several simple methods to incorporate rotational brightness variations into our phase curve fits. For example we considered a Monte Carlo approach whereby each observation was re-sampled based on its measurement uncertainty and phase curves were repeatedly fit (similar to McN21, ). We also attempted to average out rotational variation by binning reduced magnitudes nightly when fitting the phase curves. However, we found that these methods did not adequately account for rotational variations in the lightcurves. When resampling observations for objects displaying clear lightcurve variations the photometric uncertainty is generally small compared to the amplitude and so the fitted phase curve parameters and their uncertainties did not differ significantly from our original method. Similarly, binning the data did not have much effect on the final parameters. Furthermore, binning observations of objects with a wide possible range of rotational periods (minutes to days) compared to the ATLAS observational cadence would likely introduce a number of additional biases to our analysis. A more robust treatment would account for the unknown rotational phase and amplitude of each observation and so the true uncertainties of the phase curve parameters would be much wider but reflect better the true distribution of possible parameters (e.g. Waszczak et al. (2015); Alvarez-Candal et al. (2022a), SB21). However such methods are generally more computationally intensive to implement, therefore in this work we stuck with a simple phase curve fitting method, neglecting rotational effects as described in section 2.4. This allowed us to consider a sizeable sample of asteroids, and their individual apparitions, in a timely manner. As such, the formal fit uncertainties produced in this work should be used to compare relative goodness of fit within the dataset; they do not necessarily reflect the true uncertainty in the fitted phase curve parameters if rotational effects are strong.
@@ 4.5 Applicability to future surveys
In this work we have demonstrated that sparse wide field survey data can be exploited to investigate and broadly compare the properties of different asteroid populations. In particular, by considering the phase parameters and surface colours derived in this work one may compare and contrast different dynamical groups such as the asteroid families and Jupiter Trojans. However as discussed above, in order to make more precise assessments one must take into account the rotational variation introduced by the asteroid lightcurve, either in a statistical sense (e.g. Waszczak et al. (2015); Alvarez-Candal et al. (2022a), SB21) or with more detailed lightcurve shape modelling (e.g. Ďurech et al., 2020). Despite our simple phase curve model, we have built upon previous phase curve studies by considering the phase curve properties of individual apparitions for a large sample of over $100,000$ asteroids in the Main Asteroid Belt to Jupiter Trojan region, with data spanning a significant temporal baseline since 2017. By doing so we have been able to identify objects that are undergoing changes in observed absolute magnitude between apparitions. By excluding these objects from our science analysis of absolute magnitudes and colours we have improved the purity of our sample.
@@
The presence of the aforementioned apparition effects must be accounted for when analysing the results of wide field surveys expected to observe large numbers of objects over a long baseline. In particular, LSST will discover approximately $6\times 10^{6}$ MBAs over 10 years, with 100s of observations for most objects. In addition to moving object discovery, LSST will provide automatically derived properties of observed asteroids such as lightcurves, phase curves and colours (Collaboration et al., 2009). Given the volume and depth of the survey it will be difficult to followup the expected large numbers of faint asteroids, therefore LSST will likely be the only source of data for many of these objects. Considering the apparition effect is therefore essential in order to determine accurate absolute magnitudes, which directly effects studies regarding the size distribution of small Solar System bodies, such as the formation and evolution of planetesimals, properties of asteroid families, planetary defence, etc.
Furthermore, our approach which has involved detailed analysis of phase curve residuals offers a framework to monitor for unusual observations that might indicate an object of interest. We considered a range of statistics for the phase curve residuals, and in the vast majority of cases any rejected apparitions were due to poor observational coverage or low quality photometry. However, certain physical processes may also lead to deviations from the expected phase curve behaviour, e.g. rapid changes in viewing geometry (Jackson et al., 2022) or sudden increases in brightness due to the onset of activity (jewittAsteroidCometContinuum2022a). For example, asteroid 2008 GO98 was found to be active in July 2017 (Kokhirova et al., 2021). Our methodology produced a most unusual phase curve for this object, with an extreme value of $G=1675$ which was determined from an apparition spanning from 8th June until 9th December 2017. Future work in this area could involve identifying these outliers in phase curve parameter space and using the ATLAS forced photometry server to get cutouts to search for activity. Continuous updates of derived physical properties (such as absolute magnitude, phase curve and colours) as new observations are made will be an essential part of large surveys such as LSST. Therefore learning how to identify truly unusual objects in this process will greatly advance the study of relatively rare objects such as the active asteroids.
@@ 5 Conclusions
We have compiled a database of ATLAS asteroid photometry by matching the ephemerides of 4.5×$\times 10^{5}$ asteroids with accurate orbital elements to the detections produced by the ATLAS dophot pipeline, using rockAtlas (Young, 2018). We fitted phase curves in the $HG$, $H_{12}G_{12}$ models (Bowell et al., 1989; Penttilä et al., 2016) to the apparition in the $o$-filter with the best combination of number of observations and phase angle coverage. We then fitted all other apparitions for $H$ with these derived phase parameters fixed. This allowed us to assess any changes in $H$ during the survey and also to determine the $H_{c}-H_{o}$ surface colour of many asteroids. We used various metrics to select 106,055 asteroids with well defined phase curve fits for the final Selected ATLAS Phase-curves (SAP) dataset, which is available online.
@@
In this work we chose a phase curve model without rotational lightcurve variations to better investigate individual apparitions of a large number of asteroids. However, we recommend that future studies account for rotational effects, e.g. via likelihood models including rotational variation (for sparse data e.g. Schemel & Brown, 2021) and/or full CLI techniques (for objects with additional lightcurve rotation periods e.g. Ďurech et al., 2020). Accurate phase curve modelling with realistic uncertainties on absolute magnitude and phase curve parameters are essential when comparing the optical properties of different small body populations. This will be particularly relevant for large, sparse surveys such as LSST which will discover and characterise approximately $\times 10^{6}$ asteroids over its 10 year baseline (Collaboration et al., 2009). Many LSST objects will be too faint and/or numerous for dedicated followup therefore extracting accurate parameters from sparse data alone, and quickly highlighting interesting outliers, is a pressing issue.
Author’s note: Since submission we have been made aware of development of the sHG1G2 model (Carry et al., 2024); i.e. a “spinned” version of the $HG_{1}G_{2}$ Muinonen et al. (2010) model. This enhanced phase curve model also fits the oblateness and spin axis of an object in order to account for shape driven rotational effects. Such a model will certainly help to address the long-standing issues in determining accurate phase curve parameters from sparse data.